N_{2} COLLISIONAL BROADENING OF METHANE IN THE THZ REGION MEASURED AT THE SOLEIL SYNCHROTRON

\underline{V. BOUDON}, T. GABARD, Institut Carnot de Bourgogne, UMR 5209 CNRS-Université de Bourgogne, 9. Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; O. PIRALI, P. ROY, J.-B. BRUBACH, Ligne AILES – Synchrotron SOLEIL, L’Orme des Merisiers, F-91192 Gif-sur-Yvette, France; L. MANCERON, Laboratoire de Dynamique, Interactions et Réactivité, UMR 7075 – Université Pierre et Marie Curie – CNRS, Case 49, 4 Place Jussieu, F-75252, Paris Cedex, France; J. VANDER AUWERA, Service de Chimie Quantique et Photophysique, Université Libre de Bruxelles, CP 160/09, 50 avenue F.D. Roosevelt, B-1050 Brussels, Belgium.

Recently, we recorded the THz centrifugal distortion-induced spectrum of methane using synchrotron radiation at the SOLEIL facility. Intensities of pure rotation lines of CH_{4} were precisely measured, from which the induced dipole moment of this weak spectrum, of great interest for planetology, was accurately inferred.^{a} This study should in particular help to measure methane concentrations in Titan’s atmosphere.^{b} Here, we continue the work by presenting spectra of CH_{4}/N_{2} mixtures, again recorded with a 150 m optical path in a White cell and a Bruker IFS 125 HR FTIR spectrometer at the AILES beamline of SOLEIL. More precisely, the spectra were recorded with 5 % of CH_{4} in N_{2} at various total pressures (ca. 100, 200, 400, 600 and 800 mbar). These spectra will allow to measure collisional broadening and shift coefficients, as well as line-mixing parameters, with N_{2} as a perturber (like in Titan’s atmosphere).

aV. Boudon, O. Pirali, P. Roy, J.-B. Brubach, L. Manceron and J. Vander Auwera, J. Quant. Spectrosc. Radiat. Transfer, in press (2010)

bA. Coustenis, R. K. Achterberg, B. J. Conrath et al., Icarus 189, 35–62 (2007).